# ModHiFi: Identifying High Fidelity predictive components for Model Modification

### v1

@@ B.1.1 Main Proof
We first state a well known fact about Lipschitz functions. We then restate and prove Section˜3.2.
@@
Consider a network as defined in Section˜2.1. Let $\mathrm{N}_{\theta}=f^{L}\circ\ldots\circ f^{l}\circ f^{l-1:1}$ where $f^{l-1:1}=f^{l-1}\circ\ldots\circ f^{1}$. Under standard assumptions on the smoothness of activations [yu2018nisp], each layer $f^{l}$ is Lipschitz continuous with Lipschitz constant $L_{f}^{l}$. From Fact˜1,
@@
By taking an upper bound on the Lipschitz constants of each layer in the composition, we see that the subnetwork after layer $l$ has a Lipschitz constant of at least $C^{l}=\prod_{\ell>l}^{L}L_{f}^{\ell}$. Where, for convolution based networks,
@@ Appendix C Additional Experiments
    We elaborate on the Monte Carlo simulations of Equation˜k-MFS across multiple models, as well as the efficiency with which Subset Fidelity estimates this while being robust to data samples. We present noising and counterfactual results to demonstrate this.
@@ C.1.1 Monte Carlo Experiments
In this section, we provide additional details regarding Figure˜1 and demonstrate this behaviour across architectures.
Ideally, one would solve Equation˜k-MFS exactly to compute those sets that have optimal reconstructive ability at different sizes. However, since enumerating across subsets is a combinatorial problem, we instead approximate a solution by randomly sampling 1000 sets of the given size and compute the maximum across these samples. This will always provide a lower bound for the “true” curve.
Solving Equation˜k-MFS allows us to compute the optimal $(k,\eta)$-HiFi sets, since this captures the relation between $\eta$ and $k$, i.e. the tradeoff between sparsity and accuracy. We observe that in many layers (at least 50% of the model), across multiple models, there are sets which contain at most 20% of components but have a subset fidelity of around $0.8$. We also observe that as the difficulty of the task increases (CIFAR10 to ImageNet), fewer layers exhibit this sparsity, validating the assertion that networks trained on harder problems are less overparameterized.
Figures˜4, 5 and 6 show this for a ResNet-50 trained on CIFAR10, CIFAR100, and ImageNet. Figure˜7 shows this for a OPT-125m model using 128 samples of WikiText and 100 random subsets instead of a 1000. Due to the expensive nature of this experiment, we are forced to use only 100 random subsets for each size, leading to a noisier curve. However, it is clear to see that the general trend continues to hold for several layers, especially for the “down-projection" weight matrices, which are the focus of our pruning algorithm for LLMs.
@@ C.1.3 Robustness of the Fidelity Score
In Figures˜8 and 9 we show how different data sizes affect different layers in a ResNet50 model trained on CIFAR10 and ImageNet, respectively. Each data size is selected over 3 random seeds with error bars shown. For clarity, we show only a subset of layers and provide plots and code to generate them.
We observe that the values are stable for 0.2%, 0.5% and 2% of data selected, indicating that it is robust to number of samples selected.
@@
Figure 8: Fidelity scores for select layers of ResNet50 trained on ImageNet showing the effect of training and data set size.
@@
When investigating the effect of adding noise to the weights and its effect on accuracy and the fidelity score, we observe that adding zero mean noise of larger standard deviations, starting from 0.005 to 0.05, decreases the fidelity of components, with noisier weights behaving more like untrained models. We test this on a ResNet-50 trained on CIFAR10 and present the results in Figure˜10. Again, we present only a random subset of the layers for clarity.
@@
C.2 Constants in Theorem 3.2
In this section, we provided worst case and average case estimates of the constant $C_{l}$ in Section˜3.2.
In Figure˜11, we plot the constants obtained in the proof of Section˜3.2 in Section˜B.1 for a ResNet-50 trained on ImageNet, and observe that the values can be very large ($10^{38}$). However, it is important to note that these are worst case guarantees, and that these constants are much smaller in practice. In Figure˜12, we compute the ratio $\frac{\mathbb{E}\left[\|\bm{y}(\mathrm{X})-\bm{y}(\mathrm{X};\bm{M}^{l})\|^{2}\right]}{\sum_{c=1}^{c_{out}^{l}}\mathbb{E}\left[\left\|\bm{Y}^{l}_{c}(\mathrm{X})-\sum_{i\in C}m^{l}_{ci}\bm{A}^{l}_{ci}(\mathrm{X})\right\|^{2}\right]}$ and observe that these values are indeed much smaller (1010-$50$) for random values of $\bm{M}^{l}$ for the expected square loss.
Figure 11: Worst case estimates of constants $C_{l}$ for a ResNet-50 trained on ImageNet
Figure 12: Empirical estimates of constants $C_{l}$ for a ResNet-50 trained on ImageNet
@@ C.3 Discussion on the synthetic samples used in the experiments
We describe the synthetic datasets used in our vision experiments to simulate distributional access. Randomly selected example images are provided in Figure˜13. For NLP tasks, we use WikiText and Alpaca datasets [alpaca, merity2016pointersentinelmixturemodels] which are standard in this field.
@@ C.3.1 CIFAR5M
For experiments with the CIFAR10 dataset, we use CIFAR5M, a dataset containing 6 million synthetic CIFAR-10-like images sampled from a Diffusion model and labeled by a Big-Transfer model [nakkiran2021deepbootstrapframeworkgood], which we randomly sample 10,000 samples from each of the 10 classes to create our dataset. This dataset has an FID [FIDpaper] of 15.95 with respect to the CIFAR10 training set. This dataset is obtained from here.
@@ C.3.2 CIFAR100-DDPM
For experiments with the CIFAR100 dataset, we use CIFAR100-DDPM [10.5555/3540261.3540584], which we randomly downsample to contain 1,000 samples from each of the 100 classes. This dataset has an FID of 4.74 with respect to the CIFAR100 training set. We randomly sample 1,000 samples from each of the 100 classes to create our dataset. This dataset is obtained from here.
@@ C.3.3 Effect of Data Quality
To study the effect of data quality on the performance of our algorithm in vision tasks, we apply the pruning algorithm using synthetic datasets based on CIFAR10 generated with different FIDs. We use a diffusion model [Karras2022edm] to generate 3 datasets of differing quality by changing the number of diffusion steps (4,5, and 6). We report the results of our pruning algorithm with different quality datasets in Table˜5. We observe that higher quality data leads to an improved sparsity - accuracy tradeoff.
@@ C.4.2 Final ImageNet Pruned Model
In Figures˜14 and 15 we compare the final pruned models for ResNet-50 on ImageNet with DFPC [narshana2022dfpc]. We observe that our pruning algorithm removes more channels in later coupled channels than DFPC leading to higher gains in sparsity.
@@ C.4.3 Baseline selection for LLM Pruning
We choose ShortGPT [men2024shortgpt] and SliceGPT [ashkboos2024slicegpt] as baselines against which we compare ModHiFi. We do so for two broad reasons: all three methods together represent three different granularities for conducting structured pruning for LLMs, and both ShortGPT and SliceGPT are the state-of-the-art within their respective lanes.
@@ C.5 Additional Unlearning Experiments
We report additional experiments in Table˜9 on class unlearning on different architectures. For VGG-19 networks, we remove the HiFi channels for the forget class of the last 12 convolution layers. We also compare our work with DisCEdit-U from [murti2024discedit] wherein we remove discriminative components from the last 8 convolutional layers. We use a custom implementation of the algorithm for our VGG19 and ResNet50 models for CIFAR10, as those models are unavailable in the codebase of [murti2024discedit].
@@
We show that our unlearning method achieves similar or superior performance to that of [murti2024discedit] without fine-tuning. Moreover, unlike [murti2024discedit], our approach uses only synthetic samples, showing the efficacy of our work in classwise unlearning, even in the absence of training data.
@@
In this section, we compare our method when allowing for the same amount of finetuning as [jia2023model], with both synthetic data and training data access. While this violates our assumptions about loss function and training data access, we present these results to provide a fair comparison of our algorithm when run within the same constraints as our baselines. Our results can be found in Table˜12 for the Swin Transformer.
@@ C.6 Compute Platform
We implement our proposed methods in PyTorch [paszke2019pytorchimperativestylehighperformance] and use Huggingface’s transformers [wolf2020huggingfacestransformersstateoftheartnatural] for LLM implementations.
@@
We follow the inference time measurement setting of [shen2021halp, narshana2022dfpc]. Inference time is the time taken for a model to compute the forward pass for an input and does not account for loading data into memory. We compute the inference time for a batch of 640 random tensors for GPU and 64 for CPU. 100 iterations are used for warm up, after which the inference time is averaged over the next 1000 forward passes. We compute CPU and GPU measurements on a machine whose specifications can be found in Section˜C.6.
@@
We present inference time numbers with JIT compilation on Pytorch [NEURIPS2019_bdbca288].
@@ C.7.2 Training procedure
For ImageNet, we use off-the-shelf pretrained models from Torchvision [NEURIPS2019_bdbca288]. We train the model for 3 epochs after each iteration of pruning with learning rates of 0.1, 0.01, 0.001. After the pruning ends, we finally train the network for 160 epochs with a batch size of 512. We use the SGD Optimizer with a momentum factor of 0.9 and weight decay of $1\times 10^{-4}$ and start with an LR warm-up for 10 epochs, followed by Cosine Annealed step sizes with an initial learning rate of 0.1 with Cutmix and Mixup augmentations.
@@ D.1 Clarification on Lipschitz Bounds and Their Role
Empirically, we find that Subset Fidelity correlates strongly with the effect of component removal on prediction quality (see Figures˜1 and C), even in the absence of explicit Lipschitz bound estimation. This supports our design choice to treat Section˜3.2 as a motivating principle, not an operational tool.
@@ D.2 Additional Algorithmic details on Fidelity Estimation
In networks that include BatchNorm (e.g., ResNets [he2016deep], VGG [simonyan2015very]), we refine this reconstruction by centering the activations using the BatchNorm’s stored running mean. This leads to a modified formulation of the component similarity matrix:
@@
These quantities are computed efficiently using modern GPU architectures. The forward activations from a calibration set are batched and evaluated across multiple GPUs in parallel. In sequence-based architectures such as Transformers [liu2021swintransformerhierarchicalvision], we compute the expectation over all elements in the sequence dimension.
@@ D.3 Computational cost
Let $N$ be the number of data points used to estimate the saliency and $M^{l}$ be the complexity of computing the input contribution at layer $l$ for a single sample in a set of coupled channels with $m$ layers. The complexity to compute the set of retained channels for an output channel of a layer is, $t_{sal}^{l}=O(NM^{l}C_{in}^{l}d^{l})$. To select the components for the coupled channels, the top $p$ elements for each layer and output channel in them are collected, this costs $O(\sum_{l=1}^{m}C_{out}^{l}(C_{in}^{l}\log{C_{in}^{l}}+t_{sal}^{l}))$. The algorithm shows a linear dependence on the number of layers in the network, compared with the BGSC algorithm [narshana2022dfpc] which has a quadratic dependence.

### v2

@@ B.1.1 Main Proof
We first state a well-known fact about Lipschitz functions. We then restate and prove Section˜3.2.
@@
Consider a network as defined in Section˜2.1. Let $\mathrm{N}_{\theta}=f^{L}\circ\ldots\circ f^{l}\circ f^{l-1:1}$ where $f^{l-1:1}=f^{l-1}\circ\ldots\circ f^{1}$. Under standard assumptions on the smoothness of activations [94], each layer $f^{l}$ is Lipschitz continuous with Lipschitz constant $L_{f}^{l}$. From Fact˜1,
@@
By taking an upper bound on the Lipschitz constants of each layer in the composition, we see that the subnetwork after layer $l$ has a Lipschitz constant of at least $C^{l}=\prod_{\ell>l}^{L}L_{f}^{\ell}$. Where, for convolution-based networks,
@@ Appendix C Additional Experiments
    We elaborate on the Monte Carlo simulations of Eq.˜k-MFS across multiple models, as well as the efficiency with which Subset Fidelity estimates this while being robust to data samples. We present noising and counterfactual results to demonstrate this.
@@ C.1.1 Monte Carlo Experiments
In this section, we provide additional details regarding Fig.˜1 and demonstrate this behaviour across architectures.
Ideally, one would solve Eq.˜k-MFS exactly to compute those sets that have optimal reconstructive ability at different sizes. However, since enumerating across subsets is a combinatorial problem, we instead approximate a solution by randomly sampling 1000 sets of the given size and compute the maximum across these samples. This will always provide a lower bound for the “true” curve.
Solving Eq.˜k-MFS allows us to compute the optimal $(k,\eta)$-HiFi sets, since this captures the relation between $\eta$ and $k$, i.e. the tradeoff between sparsity and accuracy. We observe that in many layers (at least 50% of the model), across multiple models, there are sets which contain at most 20% of components but have a subset fidelity of around $0.8$. We also observe that as the difficulty of the task increases (CIFAR10 to ImageNet), fewer layers exhibit this sparsity, validating the assertion that networks trained on harder problems are less overparameterized.
Figs.˜4, 5 and 6 show this for a ResNet-50 trained on CIFAR10, CIFAR100, and ImageNet. Fig.˜7 shows this for a OPT-125m model using 128 samples of WikiText and 100 random subsets instead of a 1000. Due to the expensive nature of this experiment, we are forced to use only 100 random subsets for each size, leading to a noisier curve. However, it is clear to see that the general trend continues to hold for several layers, especially for the “down-projection" weight matrices, which are the focus of our pruning algorithm for LLMs.
@@ C.1.3 Robustness of the Fidelity Score
In Figs.˜8 and 9 we show how different data sizes affect different layers in a ResNet50 model trained on CIFAR10 and ImageNet, respectively. Each data size is selected over 3 random seeds, with error bars shown. For clarity, we show only a subset of layers and provide plots and code to generate them.
We observe that the values remain stable for 0.2%, 0.5%, and 2% of the data selected, indicating that the model is robust to the number of samples selected.
@@
Figure 8: Fidelity scores for select layers of ResNet50 trained on ImageNet showing the effect of training and data set size (ds).
@@
When investigating the effect of adding noise to the weights and its effect on accuracy and the fidelity score, we observe that adding zero mean noise of larger standard deviations, starting from 0.005 to 0.05, decreases the fidelity of components, with noisier weights behaving more like untrained models. We test this on a ResNet-50 trained on CIFAR10 and present the results in Fig.˜10. Again, we present only a random subset of the layers for clarity.
@@
C.2 Constants in Section˜3.2
In this section, we provided worst case and average case estimates of the constant $C_{l}$ in Section˜3.2. In Fig.˜11(a), we plot the constants obtained in the proof of Section˜3.2 in Section˜B.1 for a ResNet-50 trained on ImageNet, and observe that the values can be very large (38 orders of magnitude). However, it is important to note that these are worst case guarantees, and that these constants are much smaller in practice. In Fig.˜11(b), we compute the ratio between the global error, $\mathbb{E}\left[\|\bm{y}(\mathrm{X})-\bm{y}(\mathrm{X};\bm{M}^{l})\|^{2}\right]$ and the local error, $\sum_{c=1}^{c_{out}^{l}}\mathbb{E}\left[\left\|\bm{Y}^{l}_{c}(\mathrm{X})-\sum_{i\in C}m^{l}_{ci}\bm{A}^{l}_{ci}(\mathrm{X})\right\|^{2}\right]$ and observe that these values are indeed much smaller (1010-$50$) for random values of $\bm{M}^{l}$ for the expected square loss.
(a) Worst case
(b) Empirical
Figure 11: Estimates of constants $C_{l}$ across layers for a ResNet-50 trained on ImageNet.
@@ C.3 Discussion on the synthetic samples used in the experiments
We describe the synthetic datasets used in our vision experiments to simulate distributional access. Randomly selected example images are provided in Fig.˜12. For NLP tasks, we use WikiText and Alpaca datasets [78, 50] which are standard in this field.
@@ C.3.1 CIFAR5M
For experiments with the CIFAR10 dataset, we use CIFAR5M, a dataset containing 6 million synthetic CIFAR-10-like images sampled from a Diffusion model and labeled by a Big-Transfer model [57], which we randomly sample 10,000 samples from each of the 10 classes to create our dataset. This dataset has an FID [27] of 15.95 with respect to the CIFAR10 training set. This dataset is obtained from here.
@@ C.3.2 CIFAR100-DDPM
For experiments with the CIFAR100 dataset, we use CIFAR100-DDPM [23], which we randomly downsample to contain 1,000 samples from each of the 100 classes. This dataset has an FID of 4.74 with respect to the CIFAR100 training set. We randomly sample 1,000 samples from each of the 100 classes to create our dataset. This dataset is obtained from here.
@@ C.3.3 Effect of Data Quality
To study the effect of data quality on the performance of our algorithm in vision tasks, we apply the pruning algorithm using synthetic datasets based on CIFAR10 generated with different FIDs. We use a diffusion model [34] to generate 3 datasets of differing quality by changing the number of diffusion steps (4,5, and 6). We report the results of our pruning algorithm with different quality datasets in Table˜5. We observe that higher quality data leads to an improved sparsity - accuracy tradeoff.
@@ C.4.2 Final ImageNet Pruned Model
In Figs.˜13 and 14 we compare the final pruned models for ResNet-50 on ImageNet with DFPC [58]. We observe that our pruning algorithm removes more channels in later coupled channels than DFPC leading to higher gains in sparsity.
@@ C.4.3 Baseline selection for LLM Pruning
We choose ShortGPT [49] and SliceGPT [2] as baselines against which we compare ModHiFi. We do so for two broad reasons: all three methods together represent three different granularities for conducting structured pruning for LLMs, and both ShortGPT and SliceGPT are the state-of-the-art within their respective lanes.
@@ C.5 Additional Unlearning Experiments
We report additional experiments in Table˜9 on class unlearning on different architectures. For VGG-19 networks, we remove the HiFi channels for the forget class of the last 12 convolution layers. We also compare our work with DisCEdit-U from [55] wherein we remove discriminative components from the last 8 convolutional layers. We use a custom implementation of the algorithm for our VGG19 and ResNet50 models for CIFAR10, as those models are unavailable in the codebase of [55].
@@
We show that our unlearning method achieves similar or superior performance to that of [55] without fine-tuning. Moreover, unlike [55], our approach uses only synthetic samples, showing the efficacy of our work in classwise unlearning, even in the absence of training data.
@@
In this section, we compare our method when allowing for the same amount of finetuning as [33], with both synthetic data and training data access. While this violates our assumptions about loss function and training data access, we present these results to provide a fair comparison of our algorithm when run within the same constraints as our baselines. Our results can be found in Table˜12 for the Swin Transformer.
@@ C.6 Compute Platform
We implement our proposed methods in PyTorch [60] and use Huggingface’s transformers [87] for LLM implementations.
@@
We follow the inference time measurement setting of [70, 58]. Inference time is the time taken for a model to compute the forward pass for an input and does not account for loading data into memory. We compute the inference time for a batch of 640 random tensors for GPU and 64 for CPU. 100 iterations are used for warm up, after which the inference time is averaged over the next 1000 forward passes. We compute CPU and GPU measurements on a machine whose specifications can be found in Section˜C.6.
@@
We present inference time numbers with JIT compilation on Pytorch [61].
@@ C.7.2 Training procedure
For ImageNet, we use off-the-shelf pretrained models from Torchvision [61]. We train the model for 3 epochs after each iteration of pruning with learning rates of 0.1, 0.01, 0.001. After the pruning ends, we finally train the network for 160 epochs with a batch size of 512. We use the SGD Optimizer with a momentum factor of 0.9 and weight decay of $1\times 10^{-4}$ and start with an LR warm-up for 10 epochs, followed by Cosine Annealed step sizes with an initial learning rate of 0.1 with Cutmix and Mixup augmentations.
@@ D.1 Clarification on Lipschitz Bounds and Their Role
Empirically, we find that Subset Fidelity correlates strongly with the effect of component removal on prediction quality (see Figs.˜1 and C), even in the absence of explicit Lipschitz bound estimation. This supports our design choice to treat Section˜3.2 as a motivating principle, not an operational tool.
@@ D.2 Additional Algorithmic details on Fidelity Estimation
In networks that include BatchNorm (e.g., ResNets [26], VGG [73]), we refine this reconstruction by centering the activations using the BatchNorm’s stored running mean. This leads to a modified formulation of the component similarity matrix:
@@
These quantities are computed efficiently using modern GPU architectures. The forward activations from a calibration set are batched and evaluated across multiple GPUs in parallel. In sequence-based architectures such as Transformers [46], we compute the expectation over all elements in the sequence dimension.
@@ D.3 Computational cost
Let $N$ be the number of data points used to estimate the saliency and $M^{l}$ be the complexity of computing the input contribution at layer $l$ for a single sample in a set of coupled channels with $m$ layers. The complexity to compute the set of retained channels for an output channel of a layer is, $t_{sal}^{l}=O(NM^{l}C_{in}^{l}d^{l})$. To select the components for the coupled channels, the top $p$ elements for each layer and output channel in them are collected, this costs $O(\sum_{l=1}^{m}C_{out}^{l}(C_{in}^{l}\log{C_{in}^{l}}+t_{sal}^{l}))$. The algorithm shows a linear dependence on the number of layers in the network, compared with the BGSC algorithm [58] which has a quadratic dependence.